A fuel mixture consisting of hydrazine, N2H4, and dinitrogen tetroxide, N204,was used to launch a lunar module. These two compounds react to form nitrogen gas and water vapour. If 50.0g of hydrazine reacts with sufficient dinitrogen tetroxide, what mass of nitrogen gas is formed?
1
Expert's answer
2021-10-12T07:48:31-0400

2N2H4 + N2O4 → 3N2 + 4H2O

M(N2H4) = 32.04 g/mol

n(N2H4) "= \frac{50.0}{32.04} = 1.56 \;mol"

According to the reaction equation:

n(N2) "= \frac{3}{2}n(N_2H_4) = \frac{3}{2} \times 1.56 = 2.34 \;mol"

M(N2) = 28.01 g/mol

m(N2) "= 2.34 \times 28.01 = 65.56 \;g"

Answer: 65.56 g